# Connection formulae in the Collision Limit I: Case Studies in Lifshitz Geometry

### v2

@@ 1 Introduction
and they are usually classified into regular singularities (with $k=2$) and irregular ones (with $k>2$). Correspondingly 4d $\mathcal{N}=2$ supersymmetric gauge theories have class ${\cal S}$ construction in the context of string theory Gaiotto:2009we . In this construction, one compactifies 6d $\mathcal{N}=(2,0)$ theory on a punctured Riemann surface $\Sigma$ to obtain 4d supersymmetric gauge theories. The Riemann surface $\Sigma$ gives the well-known Seiberg-Wittern curve Seiberg:1994rs ; Seiberg:1994aj , which describes the low-energy effective theory of 4d $\mathcal{N}=2$ gauge theory. By further considering the Nekrasov-Shatashvili limit of such gauge theories Nekrasov:2009rc , the Seiberg-Witten curve is then quantized to a linear differential equation. Then a one-to-one correspondence can be established between the singularities in the linear differential equation and the punctures on $\Sigma$. That is, given a second-order linear differential equation (1), one can always find a corresponding $\mathcal{N}=2$ theory.
@@
$\psi_{\theta}^{(z_{0})}(z)=\sum_{\theta^{\prime}=\pm}A_{\theta,\theta^{\prime}}\psi_{\theta}^{(z_{1})}(z),\quad\theta=\pm,$ (6)
@@
In general, however, it is very difficult to derive the connection formulae, except for several simple examples such as the hypergeometric and Bessel functions, which have integral representations that allow one to write down such formulae. When all the singularities in the differential equation are regular (i.e. $k=2$), the linear differential equation can be obtained from the so-called BPZ equation satisfied by the correlation function in 2d CFT,
$\left(b^{-1}\frac{\partial^{2}}{\partial z^{2}}+\sum_{i=1}^{n}\frac{\partial_{t_{i}}}{z-z_{i}}+\frac{\Delta_{i}}{(z-z_{i})^{2}}\right)\langle V_{2,1}(z)\prod_{i=1}^{n}V_{i}(z_{i})\rangle=0,$ (8)
in the semiclassical limit with the central charge $c\to\infty$. We note that this fact is consistent with the Alday-Gaiotto-Tachikawa relation Alday:2009aq ; Wyllard:2009hg between 4d $\mathcal{N}=2$ theories and 2d CFTs, which states that the correlation functions in 2d CFT equal to the Nekrasov partition function of the dual gauge theory. In the case with only regular singularities, the connection formulae can be speculated from the crossing symmetry of the correlation function Bonelli:2022ten (see also Jeong:2018qpc ; Lei:2023mqx ; Liu-Zhu ). When irregular singularities appear in the linear differential equation, the situation becomes more complicated and it is even more difficult to obtain the connection formulae. In the case of one irregular singularity (with $k=3$ or $k=4$) plus one or more regular singularities, the connection formulae can be obtained with the help of the known connection formulae of the Bessel and Whittaker function and the crossing symmetry of the corresponding irregular conformal blocks Bonelli:2022ten . In the existence of the irregular singularity of $k>4$, no result has been reported in the literature. In the context of the gauge theories, for gauge theories with Lagrangian description (and thus with clear known expressions of instanton partition function), the corresponding connection formulae are rather clear, while it suddenly becomes much harder to obtain the connection formulae once we turn to the so-called non-Lagrangian theories, such as the Argyres-Douglas (AD) theories. The AD theories are theories with a strongly coupled nature Argyres:1995jj ; Xie:2012hs , and therefore there is still no known closed-form expression for the Nekrasov partition function, which causes difficulties in analysis.
In this article, we try to explore a naive idea: to directly take the collision limit of the connection formulae. The collision limit refers to the limit that brings two or more regular singularities to coincide, for example in the case of two singularities colliding into one, we have
@@ 2 Quasinormal modes in Lifshitz branes
with $r_{h}$ denoting the location of the horizon and $r=\infty$ corresponding to the boundary of the spacetime. The tortoise coordinate is given by
@@
Then we see that the QNMs can be interpreted as some of the poles in the Green function. This picture will be more convenient for our discussion when we take the collision limit, $r_{h}\rightarrow 0$.
@@
$\left(\frac{d^{2}}{dz^{2}}+\left(\frac{\gamma}{z}+\frac{\delta}{z-1}+\frac{\epsilon}{z-t}\right)\frac{d}{dz}+\frac{\alpha\beta z-q}{z(z-1)(z-t)}\right)\tilde{\psi}(z)=0,$ (33)
@@
$\tilde{\psi}_{+}^{(0)}=$ $z^{1-\gamma}\text{HeunG}[t,q-(\gamma-1)(\delta t+\epsilon),\alpha-\gamma+1,\beta-\gamma+1,2-\gamma,\delta,x],$
@@
${\rm HeunG}[t,q,\alpha,\beta,\gamma,\delta;z]=1+\frac{q}{t\gamma}z-\frac{\alpha\beta-\frac{q}{t\gamma}(\epsilon+q+\gamma+t(\gamma+\delta))}{2t(1+\gamma)}z^{2}+{\cal O}(z^{3}).$ (42)
Around each singularity, one can solve the differential equation to obtain two solutions in terms of the ${\rm HeunG}$-function, and different sets of solutions are connected by the connection formulae. For example, the solutions around $z=\infty$ are given by
$\tilde{\psi}^{(\infty)}_{+}=z^{-\alpha}{\rm HeunG}(t,q-\alpha\beta(1+t)+\alpha(\delta+t\epsilon),\alpha,\alpha-\gamma+1,$ (43)
$\alpha-\beta+1,\alpha+\beta+1-\gamma-\delta;t/z),$ (44)
@@
$\tilde{\psi}^{(\infty)}_{-}=z^{-\beta}{\rm HeunG}(t,q-\alpha\beta(1+t)+\alpha(\delta+t\epsilon),\beta,\beta-\gamma+1,$ (45)
$-\alpha+\beta+1,\alpha+\beta+1-\gamma-\delta;t/z).$ (46)
The connection formulae between $z=0$ and $z=\infty$ read
where $F$ denotes the conformal block in the semiclassical limit and $a$ is an additional parameter that should be computed from the Matone relation (154) (refer to Appendix A).
The main goal of this article is to perform consistency checks that the collision limit $r_{h}$, in which three regular singularities collide into one irregular singularity, might be taken directly at the level of the connection formulae.
@@ 4.2 Case 2: Heun function to Whittaker function
$=y^{1+\gamma-\frac{2\alpha^{\prime}}{\epsilon^{\prime}}}\exp\left(-\frac{\epsilon^{\prime}}{2y^{2}}\right){\rm HeunB\infty}[0,\alpha^{\prime}-(1+\gamma)\epsilon^{\prime},\gamma,0,-\epsilon^{\prime};y].$ (100)
We checked the identity in the last line above that relates the $\widetilde{\rm HeunB\infty}$ function with the ${\rm HeunB\infty}$ function up to $y^{6}$-order.
@@ 6 Conclusion and discussions
In this article, we performed case studies on the collision behavior $r_{h}\to 0$ of the connection formulae from the Lifshitz brane to the Lifshitz geometry. In particular, we considered two special cases: (i) $d=z+1$ (with $k=0$); (ii) $d=1$, $z=2$. In case (i), the collision limit takes the hypergeometric function to the Bessel functions, while in case (ii), the Heun functions are expected to reduce to the Whittaker function in the limit. We tested the naive idea of whether one can directly take the collision limit at the level of the connection formulae. In case (i), it works in a clear way. In case (ii), due to the complexity of the computation, we still cannot fully check it, but we provided a piece of supportive evidence that such a limit of the connection formula can be properly taken if we are given more information on the instanton partition function in the collision limit.

### v3

@@ 1 Introduction
and they are usually classified into regular singularities (with $k\leq 2$) and irregular ones (with $k>2$). Correspondingly 4d $\mathcal{N}=2$ supersymmetric gauge theories have class ${\cal S}$ construction in the context of string theory Gaiotto:2009we . In this construction, one compactifies 6d $\mathcal{N}=(2,0)$ theory on a punctured Riemann surface $\Sigma$ to obtain 4d supersymmetric gauge theories. The Riemann surface $\Sigma$ gives the well-known Seiberg-Wittern curve Seiberg:1994rs ; Seiberg:1994aj , which describes the low-energy effective theory of 4d $\mathcal{N}=2$ gauge theory. By further considering the Nekrasov-Shatashvili limit of such gauge theories Nekrasov:2009rc , the Seiberg-Witten curve is then quantized to a linear differential equation. Then a one-to-one correspondence can be established between the singularities in the linear differential equation and the punctures on $\Sigma$. That is, given a second-order linear differential equation (1), one can always find a corresponding $\mathcal{N}=2$ theory.
@@
$\psi_{\theta}^{(z_{0})}(z)=\sum_{\theta^{\prime}=\pm}A_{\theta,\theta^{\prime}}\psi_{\theta^{\prime}}^{(z_{1})}(z),\quad\theta=\pm,$ (6)
@@
In general, however, it is very difficult to derive the connection formulae, except for several simple examples such as the hypergeometric and Bessel functions, which have integral representations that allow one to write down such formulae. When all the singularities in the differential equation are regular (i.e. $k\leq 2$), the linear differential equation can be obtained from the so-called Belavin–Polyakov–Zamolodchikov (BPZ) equation satisfied by the correlation function in 2d CFT,
$\left(b^{-1}\frac{\partial^{2}}{\partial z^{2}}+\sum_{i=1}^{n}\frac{\partial_{z_{i}}}{z-z_{i}}+\frac{\Delta_{i}}{(z-z_{i})^{2}}\right)\langle V_{2,1}(z)\prod_{i=1}^{n}V_{i}(z_{i})\rangle=0,$ (8)
in the semiclassical limit with the central charge $c\to\infty$. We note that this fact is consistent with the Alday-Gaiotto-Tachikawa relation Alday:2009aq ; Wyllard:2009hg between 4d $\mathcal{N}=2$ theories and 2d CFTs, which states that the correlation functions in 2d CFT equal to the Nekrasov partition function of the dual gauge theory. In the case with only regular singularities, the connection formulae can be guessed from the crossing symmetry of the correlation function Bonelli:2022ten (see also Jeong:2018qpc ; Lei:2023mqx ; Liu-Zhu ). When irregular singularities appear in the linear differential equation, the situation becomes more complicated and it is even more difficult to obtain the connection formulae. In the case of one irregular singularity (with $k=3$ or $k=4$) plus one or more regular singularities, the connection formulae can be obtained with the help of the known connection formulae of the Bessel and Whittaker function and the crossing symmetry of the corresponding irregular conformal blocks Bonelli:2022ten . In the existence of the irregular singularity of $k>4$, no result has been reported in the literature. In the context of the gauge theories, for gauge theories with Lagrangian description (and thus with clear known expressions of instanton partition function), the corresponding connection formulae are rather clear, while it suddenly becomes much harder to obtain the connection formulae once we turn to the so-called non-Lagrangian theories, such as the Argyres-Douglas (AD) theories. The AD theories are theories with a strongly coupled nature Argyres:1995jj ; Xie:2012hs , and therefore there is still no known closed-form expression for the Nekrasov partition function, which causes difficulties in analysis.
In this article, we try to explore a naive idea: to obtain the connection formulae with irregular singularities directly by taking the collision limit of the connection formulae with only regular singularities. The collision limit refers to the limit that brings two or more regular singularities to coincide, for example in the case of two regular singularities colliding into an irregular one, we have
@@ 2 Quasinormal modes in Lifshitz branes
with $r_{h}$ denoting the location of the horizon and $r=\infty$ corresponding to the boundary (i.e. spatial infinity) of the spacetime. The tortoise coordinate is given by
@@
Remark:
We note that the asymptotic behavior of the solution $\phi$ at the spatial infinity is completely different in the Lifshitz brane (including AdS blackholes with $z=1$) and in the asymptotic flat spacetime. In the latter case, the potential $V(r)$ vanishes at the spatial infinity and thus two solutions behave as ingoing and outgoing waves, $\phi\sim e^{\pm i\omega r_{\ast}}$. However in the Lifshitz brane case, the potential (16) diverges at $r=\infty$, and two independent solutions behaves asymptotically as $\tilde{\phi}\sim r^{\frac{\pm\sqrt{z^{2}+4}-z}{2}}$. The physical solution in this case thus satisfies the Dirichlet boundary condition instead at the boundary.
$\Box$
@@
Then we see that the QNMs can be interpreted as some of the poles in the Green function (note that there are also additional poles in the variables e.g. $m_{z}$ in the Green function). This picture will be more convenient for our discussion when we take the collision limit, $r_{h}\rightarrow 0$.
@@
$\left(\frac{{\rm d}^{2}}{{\rm d}x^{2}}+\left(\frac{\gamma}{x}+\frac{\delta}{x-1}+\frac{\epsilon}{x-t}\right)\frac{\rm d}{{\rm d}x}+\frac{\alpha\beta x-q}{x(x-1)(x-t)}\right)\tilde{\psi}(x)=0,$ (33)
@@
$\tilde{\psi}_{+}^{(0)}=$ $x^{1-\gamma}\text{HeunG}[t,q-(\gamma-1)(\delta t+\epsilon),\alpha-\gamma+1,\beta-\gamma+1,2-\gamma,\delta,x],$
@@
${\rm HeunG}[t,q,\alpha,\beta,\gamma,\delta;x]=1+\frac{q}{t\gamma}x-\frac{\alpha\beta-\frac{q}{t\gamma}(\epsilon+q+\gamma+t(\gamma+\delta))}{2t(1+\gamma)}x^{2}+{\cal O}(x^{3}).$ (42)
Around each singularity, one can solve the differential equation to obtain two solutions in terms of the ${\rm HeunG}$-function, and different sets of solutions are connected by the connection formulae. For example, the solutions around $x=\infty$ are given by
$\tilde{\psi}^{(\infty)}_{+}=x^{-\alpha}{\rm HeunG}(t,q-\alpha\beta(1+t)+\alpha(\delta+t\epsilon),\alpha,\alpha-\gamma+1,$ (43)
$\alpha-\beta+1,\alpha+\beta+1-\gamma-\delta;t/x),$ (44)
@@
$\tilde{\psi}^{(\infty)}_{-}=x^{-\beta}{\rm HeunG}(t,q-\alpha\beta(1+t)+\alpha(\delta+t\epsilon),\beta,\beta-\gamma+1,$ (45)
$-\alpha+\beta+1,\alpha+\beta+1-\gamma-\delta;t/x).$ (46)
The connection formulae between $x=0$ and $x=\infty$ read
$\tilde{\psi}^{(0)}_{\theta}(x)=\sum_{\theta^{\prime}\pm}\left(\sum_{\sigma=\pm}\frac{e^{\frac{i\pi}{2}(\delta+\epsilon)}t^{\frac{\epsilon+\theta(1-\gamma)}{2}-\sigma a}e^{\frac{\theta}{2}\partial_{a_{0}}F-\frac{\theta^{\prime}}{2}\partial_{a_{\infty}}F-\frac{\sigma}{2}\partial_{a}F}}{\Gamma\left(1-\sigma a+\frac{\theta}{2}(1-\gamma)-\frac{\epsilon}{2}\right)\Gamma\left(-\sigma a+\frac{\theta}{2}(1-\gamma)+\frac{\epsilon}{2}\right)}\right.$ (47)
×$\left.\times\frac{\Gamma(1-2\sigma a)\Gamma(-2\sigma a)\Gamma(1+\theta(1-\gamma))\Gamma(-\theta^{\prime}(\alpha-\beta))}{\Gamma\left(1-\sigma a-\frac{\theta^{\prime}}{2}(\alpha-\beta)-\frac{\delta}{2}\right)\Gamma\left(-\sigma a-\frac{\theta^{\prime}}{2}(\alpha-\beta)+\frac{\delta}{2}\right)}\right)\tilde{\psi}^{(\infty)}_{\theta^{\prime}}(x),$ (48)
where $F$ denotes the conformal block in the semiclassical limit and $a$ is an additional parameter that should be computed from the Matone relation (157) (refer to Appendix A). The QNMs can be computed by requiring the ratio of the coefficients $A_{-+}/A_{--}$ to vanish, where we denoted the coefficients in the connection formula as
$\tilde{\psi}^{(0)}_{\theta}=\sum_{\theta^{\prime}=\pm}A_{\theta,\theta^{\prime}}\tilde{\psi}^{(\infty)}_{\theta^{\prime}}.$ (49)
We remark that the QNMs in the cases mentioned above are all purely imaginary, which suggests the corresponding geometry to be overdamped. This is not true for the general case of Lifshitz brane. As studied numerically in Sybesma:2015oha , for $d\leq z+1$, all the QNMs seem to be imaginary but they become complex for $d>z+1$. It will be interesting to investigate such a transition in the future work at ODE with 7 regular singularities with similar connection formulae described above.
The main goal of this article is to perform consistency checks that the collision limit $r_{h}\to 0$, in which three regular singularities collide into one irregular singularity, might be taken directly at the level of the connection formulae.
@@ 4.2 Case 2: Heun function to Whittaker function
$=\exp\left(-\frac{\epsilon^{\prime}}{2y^{2}}\right){\rm HeunB\infty}[0,\alpha^{\prime}-(1+\gamma)\epsilon^{\prime},\gamma,0,-\epsilon^{\prime};y].$ (101)
We note that the identity in the last line above that relates the $\widetilde{\rm HeunB\infty}$ function with the ${\rm HeunB\infty}$ function holds for all order. One can see this by substituting $g(y)=e^{-\frac{\epsilon^{\prime}}{2y^{2}}}h(y)$ into the biconfluent Heun equation in the $y=z^{-1}$ patch with $\delta^{\prime}=\tilde{q}=0$,
$\left[\frac{{\rm d}^{2}}{{\rm d}y^{2}}+\left(\frac{2-\gamma}{y}-\frac{\epsilon^{\prime}}{y^{3}}\right)\frac{{\rm d}}{{\rm d}y}+\frac{\alpha^{\prime}}{y^{4}}\right]g(y)=0,$ (102)
to obtain the ODE for $h(y)$ as
$\left[\frac{{\rm d}^{2}}{{\rm d}y^{2}}+\left(\frac{2-\gamma}{y}+\frac{\epsilon^{\prime}}{y^{3}}\right)\frac{{\rm d}}{{\rm d}y}+\frac{\alpha^{\prime}-(1+\gamma)\epsilon^{\prime}}{y^{4}}\right]h(y)=0.$ (103)
@@ 5.2 Case 2: Heun function to Whittaker function
Remark:
We only presented the numerical results in Figure 2 up to 9-instanton level. Due to the exponential increasing of number of terms in the instanton expansion, it is exponentially more costing to compute the higher-order corrections with the improvement slowing down very quickly. The gauge-theory approach is therefore desired to be simplified by finding some analytic pattern in the semiclassical limit of the conformal block computation to obtain high precision numerical results.
@@ 6 Conclusion and discussions
In this article, we performed case studies on the collision behavior $r_{h}\to 0$ of the connection formulae from the Lifshitz brane to the Lifshitz geometry. In particular, we considered two special cases: (i) $d=z+1$ (with $k=0$); (ii) $d=1$, $z=2$, where in both cases the collision limit takes ODEs with only regular singularities to those with one irregular singularity. In case (i), the collision limit takes the hypergeometric function to the Bessel functions, while in case (ii), the Heun functions reduce to the biconfluent Heun functions, and are further expected to be expressed in terms of the Whittaker functions in the special case of Lifshitz geometry. We tested the naive idea of whether one can directly take the collision limit at the level of the connection formulae. In case (i), it works in a clear way. In case (ii), due to the complexity of the computation, we still cannot fully check it, but we provided a piece of supportive evidence by examining a key necessary condition for the idea of collision limit to work.
In general, the collision limit considered in this article bringing three regular singularities into one irregular singularity takes a Heun function into a biconfluent Heun function, and the connection formulae for the latter is unkown so far. In this article, we focused on the reproduction of the known connection formulae in the special case in which the biconfluent Heun function degenerates to the Whittaker function. It is more desired in the future works to derive the generic connection formulae of the biconfluent Heun function by dealing the collision limit properly. The biggest obstruction at the current stage is that it is still difficult to manipulate the conformal block or equivalently the instanton partition function and its derivatives in the collision limit.
@@
Last but not least, we comment that the collision limit in the context of blackholes often means the coincidence of two horizons, e.g. the inner and outer horizons, i.e. the Cauchy and event horizons of Reissner–Nordström blackholes. This corresponds to the extremal limit of charged blackholes. Once our method of taking the collision limit of the connection formulae gets established in the future work, it will help to understand the behavior of QNMs in the extremal limit.